Cholesterol levels in men 18 to 24 years of age are normally distributed with a mean of 178 and a standard deviation of 41.

What cholesterol level corresponds to the 90th percentile, the level at which treatment may be necessary?
Answer #1

cholesterol level of man aged 18 to 24 (Normal) la 178 o=41 The 90th perentile for cholesterol level Pao = 4+20=178+1.252 x 4